In 2 years Steven will be three times as old as he was twelve years ago. next year Adam will be half Steven's age (next year as well) How old is Mark today if he is 4 times as old as Adam?
Answer by Theo(13342)

m = mark's age today.
s = steven's age today.
a = adam's age today.

in 2 years steven will be 3 times as old as he was twelve years ago.
s + 2 = 3 * (s - 12)
solve for s to get:
s = 19

next year adam will be half steven's age next year.
a + 1 = 1/2 * (19 + 1)
solve for a to get:
a = 9

mark is 4 times as old today as adam is today.
m = 4 * 9 = 36

solutions are:

steven = 19
adam = 9
mark = 36

confirm by using these numbers in the original problem statement.

in 2 years steven will be 3 times as old as he was twelve years ago.
in 2 years steven will be 21
12 years ago steven was 7
21 is equal to 3 * 7.
numbers check out.

next year adam will be half steven's age that year.
steven will be 20.
adam will be 10
1/2 20 = 10
numbers check out.

mark is 4 times as old as adam.
adam is 9.
mark is 36.
36/9 = 4
number check out.

solutions look to be good.

steven is 19
adam is 9
mark is 36